### v1

@@ 3.1 Preliminary lemmas
For a network on leaf set $X$, $\bar{\mu}(v)$ is the $|X|+1$-th dimensional vector with $\delta^{-}(v)$ as the first element and the elements of $\mu(v)$ as the latter $|X|$ elements. The modified $\mu$-representation $\bar{\mu}(\mathcal{N})$ of a network $\mathcal{N}$, is the multiset of modified $\mu$-vectors of nodes in $\mathcal{N}$.
@@
For the rest of the section, let $\mathcal{N}$ be a semi-binary stack-free network on $X$, and let $a,b\in X$ be leaves of $\mathcal{N}$. We shall show encoding results regarding modified $\mu$-representations. First, we prove some helper lemmas.
@@
Let $a$ be a leaf in $\mathcal{N}$. Then, $\mu(a)$ has multiplicity 1 or 2 in $\mu(\mathcal{N})$. If #$\#\mu(a)=1$, then its parent $p_{a}$ is a tree-node with $\mu(p_{a})\neq\mu(a)$. Furthermore, $\mu(p_{a})$ is minimal in the set $\{\mu(v):\mu(v)>\mu(a),v\in V\}$. Otherwise, if #$\#\mu(a)=2$, then $p_{a}$ is a reticulation with $\mu(p_{a})=\mu(a)$.
@@
The pair $(b,a)$ is a cherry in $\mathcal{N}$ if, and only if, $\mu_{a}(v)=\mu_{b}(v)$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$.
@@
Let us assume first that the pair $(b,a)$ is a cherry in $\mathcal{N}$ with parent node $p$. Then, for each node $v\in V\setminus\{a,b\}$ the number of paths from $v$ to either $a$ or $b$ is equal to the number of paths $P_{vp}$ from $v$ to $p$, so $\mu_{a}(v)=\mu_{b}(v)$.
For the other direction, we will use a proof by contradiction. Assume $\mu_{a}(v)=\mu_{b}(v)$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$. Now assume the pair $(b,a)$ is not a cherry in $\mathcal{N}$. This means $a$ and $b$ must have different parents $p_{a}\neq p_{b}$. However, because $\mu_{a}(v)=\mu_{b}(v)$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$, we have that $\mu_{b}(p_{a})=\mu_{a}(p_{a})=1$, therefore there is a path from $p_{a}$ to $b$. This means $p_{a}$ must be an ancestor of $p_{b}$. But also $\mu_{a}(p_{b})=\mu_{b}(p_{b})=1$, therefore $p_{b}$ must also be an ancestor of $p_{a}$. In acyclic graphs two nodes cannot be ancestors of each other unless they are the same node, therefore $p_{a}=p_{b}$, but this contradicts our assumption that $(b,a)$ is not a cherry. ∎
Note that the condition on the $\mu$-vectors implies that $\mu(a)$ and $\mu(b)$ have multiplicity 1 in $\mu(\mathcal{N})$, because if for instance $\mu(b)$ has multiplicity greater than 1 in $\mu(\mathcal{N})$, then $\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$ would still contain a vector $\mu(b)$, for which $\mu_{a}(b)=0\neq 1=\mu_{b}(b)$.
@@
The pair $(b,a)$ is a reticulated cherry in $\mathcal{N}$ with $b$ the reticulation leaf if, and only if, #$\#\mu(a)=1$, #$\#\mu(b)=2$, $\mu_{b}(v)\geq\mu_{a}(v)$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$ and $\mu(\mathcal{N})$ contains a vector $\mu(p_{a})=\mu(a)+\mu(b)$.
@@
First let us assume $(b,a)$ is a reticulated cherry in $\mathcal{N}$ with $b$ the reticulation leaf. Then the parent of $a$ is a tree node $p_{a}$ and, by Observation 1, $\mu(a)$ has multiplicity 1 in the multiset. Also, the parent of $b$ is a reticulation $p_{b}$, therefore by Observation 1, $\mu(b)$ has multiplicity 2 in the multiset. Furthermore, $p_{a}$ is the parent of $p_{b}$ and thus an ancestor of $b$. Therefore, for each path $v\rightsquigarrow p_{a}$ with $v\in V\setminus\{a,b\}$, there is at least one path $v\rightsquigarrow b$ via $p_{a}$. Furthermore, the number of paths from $v$ to $a$ equals the number of paths from $v$ to $p_{a}$. This means that $\mu_{b}(v)\geq P_{vp_{a}}=\mu_{a}(v)$ for any node $v\in V\setminus\{a,b\}$. Finally, note that $\mu(p_{a})=\mu(a)+\mu(p_{b})=\mu(a)+\mu(b)$. This proves the first direction.
@@
For the second direction, we will use proof by contradiction. Let us assume, $\mu(a)$ has multiplicity 1 in the multiset, $\mu(b)$ has multiplicity 2 in the multiset, $\mu_{b}(v)\geq\mu_{a}(v)$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$ and $\mu(\mathcal{N})$ contains $\mu(p_{a})=\mu(a)+\mu(b)$. Now assume $(b,a)$ is not a reticulated cherry. Note $\mu(p_{a})>\mu(a)$ and the only $\mu$-vectors $\mu(v)$ with $\mu(v)<\mu(p_{a})$ are $\mu(a)$ and $\mu(b)$, thus $\mu(p_{a})$ is minimal in $\{\mu(v):\mu(v)>\mu(a),v\in V\}$. Then, by Observation 1, $\mu(p_{a})$ belongs to the parent of $a$. Furthermore, by Observation 1, we know $b$ has a reticulation parent $p_{b}$. Therefore the parent $p_{a}$ of $a$ is not a parent of $p_{b}$ the parent of $b$, because otherwise $(b,a)$ would be a reticulated cherry. But there must be a path from $p_{a}$ to $b$ because $\mu_{b}(p_{a})\geq\mu_{a}(p_{a})=1$. This means there must be other nodes on the path from $p_{a}$ to $p_{b}$. Let $c$ then be the child of $p_{a}$, then $\mu(p_{a})=\mu(a)+\mu(c)=\mu(a)+\mu(b)$. Subtracting $\mu(a)$ gives $\mu(c)=\mu(b)$. This means $c$ is either the leaf $b$, which is not possible, or it is $p_{b}$, which we assumed it was not, or it is some other reticulation which has a child with the same $\mu$-vector equal to $\mu(b)$. Its child cannot be the leaf $b$, because by assumption $c$ is not $p_{b}$ and its child cannot be $p_{b}$ because the network is stack-free. Therefore, its child must be a tree-node with the same $\mu$-vector as the leaf $b$, which is not itself leaf $b$ or $p_{b}$. But, then #$\#\mu(b)=3$, which contradicts our assumption. ∎
@@
Note that if $(b,a)$ is a cherry or a reticulated cherry in $\mathcal{N}$, we say it is a cherry or reticulated cherry in $\mu(\mathcal{N})$ and $\bar{\mu}(\mathcal{N})$. Furthermore, by definition, if $(b,a)$ is a reticulated cherry, it is simple if, and only if, $\bar{\mu}(\mathcal{N})$ contains $\bar{\mu}(p_{b})=\{2\}\oplus\mu(b)$, because then $\bar{\mu}_{0}(p_{b})=2$, for $p_{b}$ the parent of $b$. Otherwise, it is complex. If $(b,a)$ is a cherry or a reticulated cherry in $\mathcal{N}$ we say that $(b,a)$ is a reducible pair in $\mathcal{N}$, in $\mu(\mathcal{N})$ and in $\bar{\mu}(\mathcal{N})$.
@@
It is important to mention here that we restrict to the class of stack-free networks to be able to identify reticulated cherries in the $\mu$-representation. Without the stack-free assumption, the conditions in Lemma 2 are not sufficient to determine whether $(b,a)$ is a reticulated cherry in general. In Figure 7, two non-isomorphic networks are displayed which have the same $\mu$-representation. To see the non-isomorphism, the left network contains a reticulated cherry while the right network does not. Clearly, $\mu$-representations do not suffice in identifying reticulated cherries if the stack-free condition is not imposed. In light of this, Cardona et al. proposed a different extended $\mu$-representation by considering, for every vertex, the number of paths to reticulations, in addition to the $\mu$-vectors. While this sufficed to prove that binary orchard networks are encoded by extended $\mu$-representations, the encoding result does not easily translate to the semi-binary network case. The networks in Figure 1 form a counterexample, they have the same extended $\mu$-representation, while they are not isomorphic.
@@ 3.2 Reconstructing orchard networks
Now let us define cherry and reticulated cherry reductions in $\bar{\mu}(\mathcal{N})$. We first define operations on vectors. Recall that to remove a vector from a multiset means to lower the multiplicity by 1 to a minimum of 0. If a vector has multiplicity 0 in a multiset we say the multiset does not contain the vector. By removing element $\bar{\mu}_{b}(v)$ from the vector $\bar{\mu}(v)$ we mean to change $\bar{\mu}(v)$ to $(\bar{\mu}_{i}(v))_{i\in\{0\}\cup X\setminus\{b\}}$. By subtracting $\bar{\mu}_{a}(v)$ from $\bar{\mu}_{b}(v)$ over $\bar{\mu}(v)$, we mean the operation of changing $\bar{\mu}(v)$ to $[\bar{\mu}_{0}(v),\bar{\mu}_{a}(v),\bar{\mu}_{b}(v)-\bar{\mu}_{a}(v),\mu_{c}(v),\ldots]$.
@@
    Remove $\bar{\mu}(b)$ from $\bar{\mu}(\mathcal{N})$.
@@
    Remove $\bar{\mu}(p_{ab})=\{1\}\oplus(\mu(a)+\mu(b))$ from $\bar{\mu}(\mathcal{N})$.
@@
    For each $\bar{\mu}(v)\in\bar{\mu}(\mathcal{N})$, remove $\bar{\mu}_{b}(v)$ from $\bar{\mu}(v)$.
Note, that because $(b,a)$ is a cherry, the parent $p_{ab}$ of $a$ and $b$ will have $\bar{\mu}$-vector $\{1\}\oplus(\mu(a)+\mu(b))$ before reduction and should be suppressed when reducing $(b,a)$. Note also that none of the in-degrees of any nodes have changed. Now let $(b,a)$ be a simple reticulated cherry in $\bar{\mu}(\mathcal{N})$, we define the simple reticulated cherry reduction of $(b,a)$ as the following operations:
@@
    Remove $\bar{\mu}(p_{a})=\{1\}\oplus(\mu(a)+\mu(b))$ from $\bar{\mu}(\mathcal{N})$.
@@
    Remove $\bar{\mu}(p_{b})=\{2\}\oplus\mu(b)$ from $\bar{\mu}(\mathcal{N})$.
@@
    For each $\bar{\mu}(v)\in\bar{\mu}(\mathcal{N})\setminus\bar{\mu}(a)$ subtract $\bar{\mu}_{a}(v)$ from $\bar{\mu}_{b}(v)$.
Note that tree-nodes have in-degree 1 and so, by Lemma 2, $\bar{\mu}(p_{a})$ is the $\bar{\mu}$-vector of the parent of $a$, which should be suppressed when reducing $(b,a)$. Furthermore, because $(b,a)$ is a simple reticulated cherry, the parent $p_{b}$ of $b$ has in-degree 2 before reducing and should be suppressed as well. Furthermore, because the edge between $p_{a}$ and $p_{b}$ there are no longer any paths to $b$ via $p_{a}$. Note that the in-degrees of any nodes that are not suppressed have not changed. Finally, we define the complex reticulated cherry reduction as follows:
@@
    Remove $\bar{\mu}(p_{a})=\{1\}\oplus(\mu(a)+\mu(b))$ from $\bar{\mu}(\mathcal{N})$.
@@
    Let $\bar{\mu}(p_{b})\in\bar{\mu}(\mathcal{N})$ be the vector with $\bar{\mu}_{0}(p_{b})>1$ and $\bar{\mu}(p_{b})=\bar{\mu}_{0}(p_{b})\oplus\mu(b)$ and lower $\bar{\mu}_{0}(p_{b})$ by 1.
@@
    For each $\bar{\mu}(v)\in\bar{\mu}(\mathcal{N})$ subtract $\bar{\mu}_{a}(v)$ from $\bar{\mu}_{b}(v)$.
For this reduction we keep the $\bar{\mu}$-vector of the parent $p_{b}$ of $b$, because it is not suppressed when $(b,a)$ is reduced in $\mathcal{N}$, because it has in-degree greater than 1 after reduction, but we do lower its in-degree by 1. Note that, by Observation 1, in stack-free networks there can only be one non-leaf node with $\mu$-vector equal to $\mu(b)$ and therefore $\bar{\mu}(p_{b})$ has multiplicity 1 in $\bar{\mu}(\mathcal{N})$. Finally note that the in-degrees of any other nodes, besides $p_{b}$ have not changed.
@@
Suppose we are given a semi-binary stack-free orchard network $\mathcal{N}_{1}$, and a semi-binary stack-free network $\mathcal{N}_{2}$ with $\bar{\mu}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{2})$. Note that $\bar{\mu}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{2})$ implies that also $\mu(\mathcal{N}_{1})=\mu(\mathcal{N}_{2})$. Then, because $\mathcal{N}_{1}$ is orchard it must contain a reducible pair of leaves $(b,a)$. If the pair $(b,a)$ is a cherry, then by Lemma 1 it must be a cherry in $\mu(\mathcal{N}_{1})$. Therefore, it is also a cherry in $\mu(\mathcal{N}_{2})$ and thus again by Lemma 1 it is a cherry in $\mathcal{N}_{2}$. In this case, if $\mathcal{N}_{1}^{\prime}$ is the network generated by reducing $(b,a)$ in $\mathcal{N}_{1}$, and $\bar{\mu}(\mathcal{N}_{1}^{\prime})$ is the $\mu$-representation of this network, then by Lemma 3, $\bar{\mu}^{\prime}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{1}^{\prime})$, where $\bar{\mu}^{\prime}(\mathcal{N}_{1})$ is the multiset generated by reducing $(b,a)$ in $\bar{\mu}(\mathcal{N}_{1})$. Thus, because there is only a single way of reducing a cherry in the $\bar{\mu}$-representation, we have that $\bar{\mu}^{\prime}(\mathcal{N}_{1})=\bar{\mu}^{\prime}(\mathcal{N}_{2})$ and by Lemma 3, $\bar{\mu}^{\prime}(\mathcal{N}_{2})$ is the $\bar{\mu}$-representation $\bar{\mu}(\mathcal{N}_{2}^{\prime})$ of the network generated by reducing $(b,a)$ in $\mathcal{N}_{2}$. To conclude, after reducing the cherry $(b,a)$ in both $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, the two networks still have the same $\bar{\mu}$-representation.
Alternatively, if $(b,a)$ is a reticulated cherry in $\mathcal{N}_{1}$ then by Lemma 2, it is a reticulated cherry in $\mu(\mathcal{N}_{1})$. Therefore, by the same argument as before, it is a reticulated cherry in $\mathcal{N}_{2}$. Furthermore, because $\bar{\mu}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{2})$, if $(b,a)$ is simple in $\mathcal{N}_{1}$ then it is simple in $\mathcal{N}_{2}$ and otherwise it is complex in both networks. As for each type of reticulated cherry there is a single way of reducing it in the $\bar{\mu}$-representation, we again see that both networks must have the same modified $\bar{\mu}$-representation after reduction of $(b,a)$.

### v2

@@ 3.1 Preliminary lemmas
For a network on leaf set $X$, $\bar{\mu}(v)$ is the $|X|+1$-th dimensional vector with $\delta^{-}(v)$ as the first coordinate and the coordinates of $\mu(v)$ as the latter $|X|$ coordinates. The modified $\mu$-representation $\bar{\mu}(\mathcal{N})$ of a network $\mathcal{N}$, is the multiset of modified $\mu$-vectors $\bar{\mu}(v)$ of nodes $v$ in $\mathcal{N}$.
@@
For the rest of the section, let $\mathcal{N}$ be a semi-binary stack-free network on $X$, and let $a,b\in X$ be leaves of $\mathcal{N}$. We shall show encoding results regarding modified $\mu$-representations. First, we prove some preliminary results.
@@
Let $a$ be a leaf in $\mathcal{N}$. Then, $\mu(a)$ has multiplicity 1 or 2 in $\mu(\mathcal{N})$. If #$\#\mu(a)=1$, then its parent $p_{a}$ is a tree-node with $\mu(p_{a})\neq\mu(a)$. Furthermore, $\mu(p_{a})$ is minimal in the multiset $\{\mu(v):\mu(v)>\mu(a),v\in V\}$. Otherwise, if #$\#\mu(a)=2$, then $p_{a}$ is a reticulation with $\mu(p_{a})=\mu(a)$.
@@
The pair $(b,a)$ is a cherry in $\mathcal{N}$ if, and only if, $\mu(v)_{a}=\mu(v)_{b}$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$.
@@
Let us assume first that the pair $(b,a)$ is a cherry in $\mathcal{N}$ with parent node $p$. Then, for each node $v\in V\setminus\{a,b\}$ the number of paths from $v$ to either $a$ or $b$ is equal to the number of paths $P_{vp}$ from $v$ to $p$, so $\mu(v)_{a}=\mu(v)_{b}$.
For the other direction, we will use a proof by contradiction. Assume $\mu(v)_{a}=\mu(v)_{b}$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$. Now assume the pair $(b,a)$ is not a cherry in $\mathcal{N}$. This means $a$ and $b$ must have different parents $p_{a}\neq p_{b}$. However, because $\mu(v)_{a}=\mu(v)_{b}$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$, we have that $\mu(p_{a})_{b}=\mu(p_{a})_{a}=1$, therefore there is a path from $p_{a}$ to $b$. This means $p_{a}$ must be an ancestor of $p_{b}$. But also $\mu(p_{b})_{a}=\mu(p_{b})_{b}=1$, therefore $p_{b}$ must also be an ancestor of $p_{a}$. In acyclic graphs two nodes cannot be ancestors of each other unless they are the same node, therefore $p_{a}=p_{b}$, but this contradicts our assumption that $(b,a)$ is not a cherry. ∎
Note that the condition on the $\mu$-vectors implies that $\mu(a)$ and $\mu(b)$ have multiplicity 1 in $\mu(\mathcal{N})$, because if for instance $\mu(b)$ has multiplicity greater than 1 in $\mu(\mathcal{N})$, then $\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$ would still contain a vector $\mu(b)$, for which $\mu(b)_{a}=0\neq 1=\mu(b)_{b}$.
@@
The pair $(b,a)$ is a reticulated cherry in $\mathcal{N}$ with $b$ the reticulation leaf if, and only if, #$\#\mu(a)=1$, #$\#\mu(b)=2$, $\mu(v)_{b}\geq\mu(v)_{a}$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$ and $\mu(\mathcal{N})$ contains a vector $\mu(p_{a})=\mu(a)+\mu(b)$.
@@
First let us assume $(b,a)$ is a reticulated cherry in $\mathcal{N}$ with $b$ the reticulation leaf. Then the parent of $a$ is a tree node $p_{a}$ and, by 1, $\mu(a)$ has multiplicity 1 in the multiset. Also, the parent of $b$ is a reticulation $p_{b}$, therefore by 1, $\mu(b)$ has multiplicity 2 in the multiset. Furthermore, $p_{a}$ is the parent of $p_{b}$ and thus an ancestor of $b$. Therefore, for each path $v\rightsquigarrow p_{a}$ with $v\in V\setminus\{a,b\}$, there is at least one path $v\rightsquigarrow b$ via $p_{a}$. Furthermore, the number of paths from $v$ to $a$ equals the number of paths from $v$ to $p_{a}$. This means that $\mu(v)_{b}\geq P_{vp_{a}}=\mu(v)_{a}$ for any node $v\in V\setminus\{a,b\}$. Finally, note that $\mu(p_{a})=\mu(a)+\mu(p_{b})=\mu(a)+\mu(b)$. This proves the first direction.
@@
For the second direction, we will use proof by contradiction. Let us assume, $\mu(a)$ has multiplicity 1 in the multiset, $\mu(b)$ has multiplicity 2 in the multiset, $\mu(v)_{b}\geq\mu(v)_{a}$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$ and $\mu(\mathcal{N})$ contains $\mu(p_{a})=\mu(a)+\mu(b)$. Now assume $(b,a)$ is not a reticulated cherry. Note $\mu(p_{a})>\mu(a)$ and the only $\mu$-vectors $\mu(v)$ with $\mu(v)<\mu(p_{a})$ are $\mu(a)$ and $\mu(b)$, thus $\mu(p_{a})$ is minimal in $\{\mu(v):\mu(v)>\mu(a),v\in V\}$. Then, by 1, $\mu(p_{a})$ belongs to the parent of $a$. Furthermore, by 1, we know $b$ has a reticulation parent $p_{b}$. Therefore the parent $p_{a}$ of $a$ is not a parent of $p_{b}$ the parent of $b$, because otherwise $(b,a)$ would be a reticulated cherry. But there must be a path from $p_{a}$ to $b$ because $\mu(p_{a})_{b}\geq\mu(p_{a})_{a}=1$. This means there must be other nodes on the path from $p_{a}$ to $p_{b}$. Let $c$ then be the child of $p_{a}$, then $\mu(p_{a})=\mu(a)+\mu(c)=\mu(a)+\mu(b)$. Subtracting $\mu(a)$ gives $\mu(c)=\mu(b)$. This means $c$ is either the leaf $b$, which is not possible, or it is $p_{b}$, which we assumed it was not, or it is some other reticulation which has a child with the same $\mu$-vector equal to $\mu(b)$. Its child cannot be the leaf $b$, because by assumption $c$ is not $p_{b}$ and its child cannot be $p_{b}$ because the network is stack-free. Therefore, its child must be a tree-node with the same $\mu$-vector as the leaf $b$, which is not itself leaf $b$ or $p_{b}$. But, then #$\#\mu(b)=3$, which contradicts our assumption. ∎
Let us now define cherries and reticulated cherries as subsets of the original and modified $\mu$-representation as follows naturally from Lemma 1 and Lemma 2.
Definition 3.
  • The pair $(b,a)$ is a cherry in $\mu(\mathcal{N})$ if $\mu(v)_{a}=\mu(v)_{b}$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$.
  • The pair $(b,a)$ is a cherry in $\bar{\mu}(\mathcal{N})$ if $\bar{\mu}(v)_{a}=\bar{\mu}(v)_{b}$ for each $\bar{\mu}(v)\in\bar{\mu}(\mathcal{N})\setminus\{\bar{\mu}(a),\bar{\mu}(b)\}$.
  • The pair $(b,a)$ is a reticulated cherry in $\mu(\mathcal{N})$ with $b$ the reticulation leaf if #$\#\mu(a)=1$, #$\#\mu(b)=2$, $\mu(v)_{b}\geq\mu(v)_{a}$ for each $\mu(v)\in\mu(\mathcal{N})\setminus\{\mu(a),\mu(b)\}$ and $\mu(\mathcal{N})$ contains a vector $\mu(p_{a})=\mu(a)+\mu(b)$.
  • The pair $(b,a)$ is a reticulated cherry in $\bar{\mu}(\mathcal{N})$ with $b$ the reticulation leaf if the following hold:
    • there does not exist $\bar{\mu}(v)\in\bar{\mu}(\mathcal{N})$ with $\bar{\mu}(v)_{i}=\bar{\mu}(a)_{i}$ for $i\in X$,
    • there exists a vector $\bar{\mu}(p_{b})\in\bar{\mu}(\mathcal{N})$ with $\bar{\mu}(v)_{i}=\bar{\mu}(b)_{i}$ for $i\in X$,
    • $\bar{\mu}(v)_{b}\geq\bar{\mu}(v)_{a}$ for each $\bar{\mu}(v)\in\bar{\mu}(\mathcal{N})\setminus\{\bar{\mu}(a),\bar{\mu}(b)\}$ and
    • $\bar{\mu}(\mathcal{N})$ contains a vector $\bar{\mu}(p_{a})$ s.t. $\bar{\mu}(p_{a})_{i}=\bar{\mu}(a)_{i}+\bar{\mu}(b)_{i}$ for $i\in X$.
    It is simple if $\bar{\mu}(\mathcal{N})$ contains $\bar{\mu}(p_{b})=[2]\oplus\mu(b)$ and otherwise it is complex.
@@
Note that if $(b,a)$ is a cherry or a reticulated cherry in $\mathcal{N}$, it is a cherry or reticulated cherry in $\mu(\mathcal{N})$ and $\bar{\mu}(\mathcal{N})$. Furthermore, if $(b,a)$ is a reticulated cherry, it is simple if, and only if, $\bar{\mu}(\mathcal{N})$ contains $\bar{\mu}(p_{b})=[2]\oplus\mu(b)$, because then $\bar{\mu}(p_{b})_{0}=2$, for $p_{b}$ the parent of $b$. Otherwise, it is complex. If $(b,a)$ is a cherry or a reticulated cherry in $\mathcal{N}$ we say that $(b,a)$ is a reducible pair in $\mathcal{N}$, in $\mu(\mathcal{N})$ and in $\bar{\mu}(\mathcal{N})$.
@@
It is important to mention here that we restrict to the class of stack-free networks to be able to identify reticulated cherries in the $\mu$-representation. Without the stack-free assumption, the conditions in 2 are not sufficient to determine whether $(b,a)$ is a reticulated cherry in general. In Figure 7, two non-isomorphic networks are displayed which have the same $\mu$-representation. To see the non-isomorphism, the left network contains a reticulated cherry while the right network does not. Clearly, $\mu$-representations do not suffice in identifying reticulated cherries if the stack-free condition is not imposed. In light of this, Cardona et al. proposed a different extended $\mu$-representation by considering, for every vertex, the number of paths to reticulations, in addition to the $\mu$-vectors. While this sufficed to prove that binary orchard networks are encoded by extended $\mu$-representations, the encoding result does not easily translate to the semi-binary network case. The networks in Figure 1 form a counterexample, they have the same extended $\mu$-representation, while they are not isomorphic.
@@ 3.2 Reconstructing orchard networks
Now let us define cherry and reticulated cherry reductions in $\bar{\mu}(\mathcal{N})$. We first define operations on vectors. Recall that to remove a vector from a multiset means to lower the multiplicity by 1 to a minimum of 0. If a vector has multiplicity 0 in a multiset we say the multiset does not contain the vector.
@@
    Remove the unit vector $\bar{\mu}(b)$ from $\bar{\mu}(\mathcal{N})$.
@@
    Remove the vector $\bar{\mu}(p_{ab})=[1]\oplus(\mu(a)+\mu(b))$ from $\bar{\mu}(\mathcal{N})$.
@@
    For each $\bar{\mu}(v)$ in $\bar{\mu}(\mathcal{N})$, replace it with a vector $(\bar{\mu}(v)_{i})_{i\in S}$ where $S$ is the set $\{0\}\cup X\setminus\{b\}$.
Note, that because $(b,a)$ is a cherry, the parent $p_{ab}$ of $a$ and $b$ will have $\bar{\mu}$-vector $[1]\oplus(\mu(a)+\mu(b))$ before reduction and should be suppressed when reducing $(b,a)$. Note also that none of the in-degrees of any nodes have changed. Now let $(b,a)$ be a simple reticulated cherry in $\bar{\mu}(\mathcal{N})$, we define the simple reticulated cherry reduction of $(b,a)$ as the following operations:
@@
    Remove the vector $\bar{\mu}(p_{a})=[1]\oplus(\mu(a)+\mu(b))$ from $\bar{\mu}(\mathcal{N})$.
@@
    Remove the vector $\bar{\mu}(p_{b})=[2]\oplus\mu(b)$ from $\bar{\mu}(\mathcal{N})$.
@@
    For each $\bar{\mu}(v)\in\bar{\mu}(\mathcal{N})\setminus\bar{\mu}(a)$ replace it with the vector $\bar{\mu}(v)-\bar{\mu}(v)_{a}([0]\oplus\mu(b))$.
Note that tree-nodes have in-degree 1 and so, by 2, $\bar{\mu}(p_{a})$ is the $\bar{\mu}$-vector of the parent of $a$, which should be suppressed when reducing $(b,a)$. Furthermore, because $(b,a)$ is a simple reticulated cherry, the parent $p_{b}$ of $b$ has in-degree 2 before reducing and should be suppressed as well. Furthermore, because the edge between $p_{a}$ and $p_{b}$ there are no longer any paths to $b$ via $p_{a}$. Note that the in-degrees of any nodes that are not suppressed have not changed. Finally, we define the complex reticulated cherry reduction as follows:
@@
    Remove the vector $\bar{\mu}(p_{a})=[1]\oplus(\mu(a)+\mu(b))$ from $\bar{\mu}(\mathcal{N})$.
@@
    Let $\bar{\mu}(p_{b})\in\bar{\mu}(\mathcal{N})$ be the vector with $\bar{\mu}(p_{b})_{0}>1$ and $\bar{\mu}(p_{b})=\bar{\mu}(p_{b})_{0}\oplus\mu(b)$ and lower $\bar{\mu}(p_{b})_{0}$ by 1.
@@
    For each $\bar{\mu}(v)\in\bar{\mu}(\mathcal{N})\setminus\bar{\mu}(a)$ replace it with the vector $\bar{\mu}(v)-\bar{\mu}(v)_{a}([0]\oplus\mu(b))$.
For this reduction we keep the $\bar{\mu}$-vector of the parent $p_{b}$ of $b$, because it is not suppressed when $(b,a)$ is reduced in $\mathcal{N}$, because it has in-degree greater than 1 after reduction, but we do lower its in-degree by 1. Note that, by 1, in stack-free networks there can only be one non-leaf node with $\mu$-vector equal to $\mu(b)$ and therefore $\bar{\mu}(p_{b})$ has multiplicity 1 in $\bar{\mu}(\mathcal{N})$. Finally note that the in-degrees of any other nodes, besides $p_{b}$ have not changed.
@@
Suppose we are given a semi-binary stack-free orchard network $\mathcal{N}_{1}$, and a semi-binary stack-free network $\mathcal{N}_{2}$ with $\bar{\mu}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{2})$. Note that $\bar{\mu}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{2})$ implies that also $\mu(\mathcal{N}_{1})=\mu(\mathcal{N}_{2})$. Then, because $\mathcal{N}_{1}$ is orchard it must contain a reducible pair of leaves $(b,a)$. If the pair $(b,a)$ is a cherry, then by 1 it must be a cherry in $\mu(\mathcal{N}_{1})$. Therefore, it is also a cherry in $\mu(\mathcal{N}_{2})$ and thus again by 1 it is a cherry in $\mathcal{N}_{2}$. In this case, if $\mathcal{N}_{1}^{\prime}$ is the network generated by reducing $(b,a)$ in $\mathcal{N}_{1}$, and $\bar{\mu}(\mathcal{N}_{1}^{\prime})$ is the $\mu$-representation of this network, then by 3, $\bar{\mu}^{\prime}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{1}^{\prime})$, where $\bar{\mu}^{\prime}(\mathcal{N}_{1})$ is the multiset generated by reducing $(b,a)$ in $\bar{\mu}(\mathcal{N}_{1})$. Thus, because there is only a single way of reducing a cherry in the $\bar{\mu}$-representation, we have that $\bar{\mu}^{\prime}(\mathcal{N}_{1})=\bar{\mu}^{\prime}(\mathcal{N}_{2})$ and by 3, $\bar{\mu}^{\prime}(\mathcal{N}_{2})$ is the $\bar{\mu}$-representation $\bar{\mu}(\mathcal{N}_{2}^{\prime})$ of the network generated by reducing $(b,a)$ in $\mathcal{N}_{2}$. To conclude, after reducing the cherry $(b,a)$ in both $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, the two networks still have the same $\bar{\mu}$-representation.
Alternatively, if $(b,a)$ is a reticulated cherry in $\mathcal{N}_{1}$ then by 2, it is a reticulated cherry in $\mu(\mathcal{N}_{1})$. Therefore, by the same argument as before, it is a reticulated cherry in $\mathcal{N}_{2}$. Furthermore, because $\bar{\mu}(\mathcal{N}_{1})=\bar{\mu}(\mathcal{N}_{2})$, if $(b,a)$ is simple in $\mathcal{N}_{1}$ then it is simple in $\mathcal{N}_{2}$ and otherwise it is complex in both networks. As for each type of reticulated cherry there is a single way of reducing it in the $\bar{\mu}$-representation, we again see that both networks must have the same modified $\bar{\mu}$-representation after reduction of $(b,a)$.